OpenStudy (anonymous):

determine the center and the radius of the circle: x^2/2+y^2/2=18

OpenStudy (bradely):

x^2+y^2=36 is in the form (x-a)^2+(y-b)^2 =r^2 center (a,b)=(0,0) Radius r= sqrt(36)=6 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

how about center at origin,circle passes through (4,1)